Please help me solve this problem with a deck of cards, It has two parts.
You separate the 12 face cards out of a deck of 52 cards.
Assume that the remaining cards have been shuffle. Select 3 cards from the pile of face cards. Show how you derive the sulution. Tell if combination or permutation and why.
Part 1. How many ways are there of selecting one of each face card from the pile?
Part 2. How many ways are there of selecting 3 of the same face cards (ie, 3 jacks, 3 queens, 3 kings).
For part 1 my answer is 64 ways. combination, reason was not in any specific order.
for part 2, I said 12 ways, combination, reason not any speific order.
My teacher said I was wrong.
Can you show me every step. Please!!

Found 2 solutions by stanbon, near2u_28@yahoo.com:
Answer by stanbon(75887)

You separate the 12 face cards out of a deck of 52 cards.
Assume that the remaining cards have been shuffle.
Select 3 cards from the pile of face cards. Show how you derive the sulution. Tell if combination or permutation and why.
-------------------------
Part 1. How many ways are there of selecting one of each face card from the pile?
# of ways to pick a king: 4C1 = 4
# of ways to pick a queen: 4C1 = 44
# of ways to pick a jack: 4C1 = 4
# of ways to pick one of each face card from the pile: 4*4*4 =64
You are applying the General counting Rule.
-------------------------------
Part 2. How many ways are there of selecting 3 of the same face cards (ie, 3 jacks, 3 queens, 3 kings).
3 jacks: 4C3 = 4
3 queens: 4C3 = 4
3 kings: 4C3 = 4
# of ways to pick 3 kings, 3 queens, 3 jacks: 4*4*4 = 64
---------------
Cheers,
Stan H.

Answer by near2u_28@yahoo.com(2)

You separate the 12 face cards out of a deck of 52 cards.
Assume that the remaining cards have been shuffle.
Select 3 cards from the pile of face cards. Show how you derive the sulution. Tell if combination or permutation and why.
-------------------------
Part 1. How many ways are there of selecting one of each face card from the pile?
# of ways to pick a king: 4C1 = 4
# of ways to pick a queen: 4C1 = 44
# of ways to pick a jack: 4C1 = 4
# of ways to pick one of each face card from the pile: 4*4*4 =64
You are applying the General counting Rule.
-------------------------------
Part 2. How many ways are there of selecting 3 of the same face cards (ie, 3 jacks, 3 queens, 3 kings).
3 jacks: 4C3 = 4
3 queens: 4C3 = 4
3 kings: 4C3 = 4
# of ways to pick 3 kings, 3 queens, 3 jacks: 4*4*4 = 64
---------------